OpenStudy (cyclone3260):

Evaluate x^-2 for x = -3

OpenStudy (solomonzelman):

well there is a formula. \[\huge\color{red}{a^{-b}=\frac{1}{a^b}}\]so using this, \[\huge\color{blue}{x^{-2}=?}\]

OpenStudy (solomonzelman):

when you get it, plug 3 for x.

OpenStudy (solomonzelman):

@Cyclone3260, do you understand this now?

OpenStudy (cyclone3260):

1 _ x^-2 <-- Something like this?

OpenStudy (cyclone3260):

1 - -6x

OpenStudy (solomonzelman):

\[\huge\color{red}{x^{-2}=\frac{1}{x^2}}\]and knowing that x=3, then, \[\huge\color{red}{\frac{1}{(3)^2}}~~~~->~~~~~\huge\color{green}{\frac{1}{9}}\]

OpenStudy (solomonzelman):

my bad you put in -3 for x, but the answer is still same.

OpenStudy (cyclone3260):

Thank you!

OpenStudy (solomonzelman):

Anytime!

OpenStudy (cyclone3260):

But what made the denominator be 9?

OpenStudy (cyclone3260):

Oh I'm sorry, I get it now.

OpenStudy (solomonzelman):

yes, it's \[\huge\color{purple}{\frac{1}{9}}\]